QUESTION 4 the magnitude of the force on a 03 С charge if there is a ОА С charge 2 m directly above t and a 0.2 С charge 5 m directly below it is--N. QUESTION 5 Themagnitude of the force on a accharge if there is a-0.4 C charge 2 m direct y above it and a 0.2 С charge Sm direct ybelow is-N. QUESTION 6 The force on a 0.3 C charge at the origin of an xy plane, if there is a 0.4 C charge at position (2.2 ts equal to the vector Round each component to the nearest newton N. NJ QUESTION 7 the force on a 0.3 C charge at the origin of an xy plane if there is a-Q4 C charge at position 2 2, and a o.e charge at positiona īzi, equal to the yector ( N. Round each component to the nearest newton QUESTION 8 is the amount of charge in the universe the same as it was 1000 years MEo? O yes Chiek Save and Submit to save and submit. Click Save All Anavera to soue all aneers
Answer #1

4.

F = k q ( q1/ r1^2 - q2 / r2^2)

F = 9*10^9* 0.3* ( 0.4/ 2^2 - 0.2/ 5^2)

F = 2.484*10^8 N

========

5.

F = 9*10^9* 0.3* ( 0.4 / 2^2 + 0.2/ 5^2)

F = 2.916*10^8 N

========

6.

F = 9*10^9* 0.3*0.4 / (2^2 + 2^2) = 1.35*10^8 N

in component form,

F = -( 1.35 cos 45 i - 1.35 cos 45 j) * 10^8

F = - 10^8 i - 10^8 j

========

Plz post other problem seperately

Comment in case any doubt, will reply for sure.. Goodluck